# Hybrid Bernstein Normalizing Flows for Flexible Multivariate Density Regression with Interpretable Marginals

### v1

@@ 1 Introduction
Many real-world regression problems involve modeling of complex conditional distributions of an outcome $Y$ given an input vector $\mathbf{x}=[x_{1},\ldots,x_{D}]^{\top}$. In such instances, a mere prediction of the mean would be insufficient to capture the aleatoric uncertainty inherent in the data-generating process and a comprehensive understanding requires estimating the conditional distribution $Y|\mathbf{x}$ [Kneib et al., 2021]. This becomes more challenging for multivariate outcomes $\mathbf{Y}\in\mathbb{R}^{D}$, both in terms of estimation and model interpretation. While normalizing flow (NF[Papamakarios et al., 2017] models can represent such complex conditional multivariate distributions, their deep neural network architectures often hinder interpretability. Statistical methods for flexible density estimation in high dimensions have largely focused on graphical models, often involving copulas [e.g., Liu et al., 2011, Bauer and Czado, 2016, Nagler and Czado, 2016]. A notable exception is the MCTM, a transformation model approach [Hothorn et al., 2018] that shares similarities with NFs and models the joint relationship similar to copulas. While less flexible than NFs, MCTMs preserve the interpretability of the feature-outcome relationship.
@@ 2 Background and Related Work
Several methods have emerged for modeling complex conditional distributions. Among these methods, transformation-based models have proven to be highly flexible and have been extensively used since their introduction by Box and Cox [1964].
@@ 2.1 Probability Transformation
Transformation models are based on the probability-transformation theorem, using a bijective and continuously differentiable transformation function $h$ to map a complex distribution $F_{Y}$ to a simpler one $F_{Z}$ (often standard normal) [Kneib et al., 2021] via $F_{Y}=F_{Z}(h(y)).$ This allows density estimation without restrictive shape assumptions [Hothorn et al., 2018]. The density $f_{Y}$ is calculated from the base density $f_{Z}$ by
@@ 2.2 Normalizing Flows and Transformation Models
NFs apply the probability-transformation theorem using a series of $K$ simple bijective transformation functions $h_{k}$ to form more expressive transformations $h(z)=h_{K}\circ h_{K-1}\circ\dots\circ h_{1}(z)$ [see, e.g. Papamakarios et al., 2021, for a comprehensive review]. Other methods use a single flexible transformation like sum-of-squares polynomials [Jaini et al., 2019] or splines [Durkan et al., 2019]. Conditional transformation models (CTMs) also focus on single transformation function, which, however, can be easily interpreted [see, e.g., Hothorn et al., 2018]. This idea has also recently been combined with neural-networks [Baumann et al., 2021, Sick et al., 2021, Arpogaus et al., 2023, Kook et al., 2024].
A fundamental difference between NFs and CTMs is the definition of the transformation direction [Kook et al., 2024]: NFs transform a simple base density $f_{Z}(z)$ into a more complex target density $f_{Y}(y)$ through a series of transformations. CTMs do it vice-versa, using only a single transformation function, often utilizing flexible Bernstein polynomials (see Section A.1 for details).
CTMs allow interpretability through the use of structured additive predictors (SAPs) [Klein et al., 2019, Rügamer et al., 2023] (see Sections A.5 and B.3). The black-box character of neural networks in NFs hinders interpretability, and variants using decision trees [Papastefanopoulos et al., 2025] and graphical models [Wehenkel and Louppe, 2021] have been explored in recent years.
@@ 2.3 Multivariate and Conditional Models
To model a potentially multivariate conditional distribution $F_{\mathbf{Y}|\mathbf{X}=\mathbf{x}}$, we define $z_{j}=h(y_{j},\bm{\theta}_{j})\text{ with }\bm{\theta}_{j}=c_{j}(\mathbf{y},\mathbf{x})$ where $h$ is a bijective transformation function for the element $y_{j}$ with parameters $\bm{\theta}_{j}$ obtained from the conditioner $c_{j}$ [Papamakarios et al., 2021]. The latter can be an arbitrarily complex function that depends on the features $\mathbf{x}$ and, in the multivariate case, on some variables of the response $\mathbf{y}$. The limiting factor here is the computational tractability of the Jacobian in Equation 1.
@@
To make computations tractable and provide a model that is better to interpret, MCTMs define a triangular transformation $H(\bm{y},\bm{\Theta}(\bm{x}))=\left(h_{1}(y_{1}|\bm{\theta}_{1}),\ldots,h_{J}(y_{1},\ldots,y_{J}|\bm{\theta}_{J})\right)^{\top}$. This Tranformation is characterized by a set of linearly combined marginal basis transformations $\tilde{h}_{j}(y_{j}|\bm{\theta}_{j})$, scaled with a $(J\times J)$ triangular coefficient matrix $\bm{\Lambda}$ to encode structural information: $h_{j}(y_{j},\ldots,y_{j}|\bm{\theta}_{j})=\lambda_{j,1}\tilde{h}_{1}(y_{1}|\bm{\theta}_{1})+\ldots+\lambda_{j,j-1}\tilde{h}_{j-1}(y_{j-1}|\bm{\theta}_{j-1})+\tilde{h}_{j}(y_{j}|\bm{\theta}_{j})$. The function $\bm{\Theta}:\bm{x}\mapsto\bm{(}\Lambda,\bm{\theta}_{1},\ldots,\bm{\theta}_{J})$ accounts for feature effects on the parameters [Klein et al., 2019]. Hence, MCTMs can only model linear dependencies between the response variables $y_{j}$, making them an inadequate choice for complex joint distributions as shown in Figure 1. This limitation motivates the development of more flexible approaches, such as the hybrid model proposed in this paper.
@@
In the NF literature, various approaches are used to deal with computations involving the aforementioned Jacobian. Among these, autoregressive models are one of the most prevalent ones. Due to the autoregressive structure, these NFs yield a triangular Jacobian and thus the determinant simplifies to the product of the Jacobian’s diagonal entries [Papamakarios et al., 2021, Kobyzev et al., 2021]. These models factorize a multivariate distribution based on the chain rule of probability. They are typically implemented by conditioning the transformation $h_{j}$ of the $j$th element of $\mathbf{y}$ on all previous elements $\mathbf{y}_{<j}=(y_{1},\dots,y_{j-1})$ or a subset of those. This can be seen as a non-linear generalization of the triangular coefficient matrix $\Lambda$ used in MCTMs [Kobyzev et al., 2021]. There are two widely adopted neural-network architectures for implementing an autoregressive conditioner $c_{j}$s [Papamakarios et al., 2017] and CFs [Dinh et al., 2017].
MAFs generalize this idea by implementing the conditioner with neural networks that incorporate autoregressive constraints through parameter masking, inspired by the masked autoencoder for distribution estimation (MADE) architecture [Germain et al., 2015]. The main advantage is that the parameters can be obtained in one neural network pass $\left(\bm{\theta}_{1},\ldots,\bm{\theta}_{J}\right)^{\top}=\left(c_{1}(),c_{2}(y_{1}),\ldots,c_{J}(\mathbf{y}_{<j})\right)^{\top}$. Further, if $h$ and the conditioning network are expressive enough, they remain universal approximators to transform between any two distributions [Papamakarios et al., 2021]. In practice, however, the results highly depend on the ordering of the input variables. Because some orderings can be extremely difficult to learn, multiple MAFs layers are often stacked with permutations between.
@@ 3.1 Model Specification
This involves optimizing the parameters of the marginal transformation functions $H_{1}$ and the parameters of the autoregressive flow $H_{2}$, which are controlled by masked neural networks. Optimization can be done using any suitable gradient-based optimization algorithm, such as Adam [Kingma and Ba, 2017]. Once the model is trained, we can sample from the joint distribution of $\mathbf{Y}|\mathbf{X}$ by:
@@ 3.2 Interpretability
Our hybrid approach combines CTM’s interpretability with the flexibility of autoregressive NFs: The marginal transformation step ($H_{1}$) may use shifted Bernstein polynomials, which allows quantifying feature effects on the marginal distribution, similar to coefficients in linear models or generalized additive models (GAMs) [Wood, 2017]. For instance, with a logistic base distribution, linear effect coefficients in the SAP can be interpreted as linear changes on the level of log-odds ratios (see Sections A.5 and B.3 for more details). Our model prioritizes marginal interpretability while accepting a trade-off for the dependence structure to gain flexibility via the masked autoregressive flow in step two ($H_{2}$). It is suitable for scenarios requiring understanding individual feature effects while modeling complex relationships between response variables. Understanding how features affect the dependence structure in this step is challenging and presents another open research question.
@@ 4 Numerical Experiments
Next, we evaluate our method using simulated and real-world datasets. All models are implemented using TensorFlow (v2.15.1) and TensorFlow Probability (v0.23.0) and trained using Adam [Martín Abadi et al., 2015, Dillon et al., 2017, Kingma et al., 2017]. In our experiments, hyperparameters were chosen based on our previous experience and settings suggested in related works. However, we did some automated hyperparameter tuning (using Optuna) on some models applied to the simulated 2D data with performance monitored on a held-out validation set. This was mainly done to understand the influence of certain parameters, but it was not necessary to obtain a good model for each data set. We found that high-order Bernstein polynomials are needed for some datasets, suggesting the fixed knot placement of Bernstein polynomials could limit their flexibility[see Hothorn et al., 2018, for an ablation to the sensitivity of the Bernstein order].
@@ 4.1 Benchmark Datasets
We evaluate our method on five common benchmark datasets POWER, GAS, HEPMASS, MINIBOONE, and BSDS300 (see Appendix Section B.5). We follow the preprocessing steps as in Papamakarios et al. [2017] and compare a MAF and our proposed hybrid MAF (HMAF). Both models utilize masked autoregressive transformations. Each MAF layer uses a masked affine transformation, parameterized by a masked neural network followed by an invertible linear transformation (1x1 convolution) initialized with a random (non-trainable) permutation. For the HMAF, we use a marginal transformation step with Bernstein polynomials before the MAFs layers.
For both MAFs and HMAF, we employ Rational Quadratic Splines (RQS) as the transformation functions within the MAFs layers of $H_{2}$. Hyperparameters, including the number of MAFs layers, the number of hidden units in the MADE networks, and the number of bins for the RQS transformations, were chosen based on the values reported in Durkan et al. [2019] and are detailed in the supplementary material and the GitHub repository. All models were trained using the Adam optimizer with early stopping after $50$ epochs without improvements and a cosine learning rate schedule.
@@ 4.2 Simulated Data
We start with the classical bivariate (i.e. $J=2$) NFs datasets, moons and circles, exhibiting non-linear dependencies. Each dataset has $16,384$ data points (we reserved 25% for validation) generated using scikit-learn [Pedregosa et al., 2011]. A binary feature $x$ based on spatial location is introduced. For moons, $x=1$ indicates the lower-right moon, and $x=0$ the upper-left. For circles, $x=1$ denotes the inner circle, and $x=0$ the outer. This assesses the models’ ability to capture $f(\mathbf{y}|x)$. We compare the following models:
@@
  • multivariate conditional transformation model (MCTM): Employs Bernstein polynomials of order $M=300$ for marginal transformations and a triangular matrix $\Lambda$ for the dependency structure. In the conditional case, the marginal shift parameter $\beta_{x}$ and $\Lambda_{x}$ are obtained using Bernstein polynomials of order 6, as in Klein et al. [2019]. While this approach offers interpretability through SAP, the linear dependence structure constrains its capacity.
@@
All models use the Adam optimizer [Kingma and Ba, 2017] with early stopping and cosine learning rate decay [Loshchilov and Hutter, 2017]. All chosen hyperparameters are reported in the supplementary material.
@@ 4.3 Malnutrition Data
Finally, we evaluate our approach using a real-world dataset on childhood malnutrition in India, with three anthropometric indices (stunting, wasting, and underweight) as response variables. The goal is to model the joint distribution conditional on the child’s age (cage). We follow the data preprocessing steps from Klein et al. [2019].
@@
  • MCTM: Identical to the model specification in Klein et al. [2019], this model combines marginal transformations with a triangular matrix $\Lambda$. We capture the influence of cage on both the marginals (using linear shifts) and on the elements of $\Lambda$ using Bernstein polynomials of order $M=6$, i.e. the marginal shift is modeled as Bernstein polynomial $\bm{\beta}_{x}=\alpha(\text{cage})^{\top}\bm{\vartheta}$, where $\alpha$ represents the Bernstein basis, and $\bm{\vartheta}$ are the coefficients.

### v2

@@ 1 Introduction
Many real-world regression problems involve modeling of complex conditional distributions of an outcome $Y$ given an input vector $\mathbf{x}=[x_{1},\ldots,x_{D}]^{\top}$. In such instances, a mere prediction of the mean would be insufficient to capture the aleatoric uncertainty inherent in the data-generating process and a comprehensive understanding requires estimating the conditional distribution $Y|\mathbf{x}$ (Kneib2021). This becomes more challenging for multivariate outcomes $\mathbf{Y}\in\mathbb{R}^{D}$, both in terms of estimation and model interpretation. While normalizing flow (NF(Papamakarios2018) models can represent such complex conditional multivariate distributions, their deep neural network architectures often hinder interpretability. Statistical methods for flexible density estimation in high dimensions have largely focused on graphical models, often involving copulas (e.g., Liu2011; Bauer2016; Nagler2016). A notable exception is the MCTM, a transformation model approach (Hothorn2018) that shares similarities with NFs and models the joint relationship similar to copulas. While less flexible than NFs, MCTMs preserve the interpretability of the feature-outcome relationship.
@@ 2 Background and Related Work
Several methods have emerged for modeling complex conditional distributions. Among these methods, transformation-based models have proven to be highly flexible and have been extensively used since their introduction by Box1964.
@@ 2.1 Probability Transformation
Transformation models are based on the probability-transformation theorem, using a bijective and continuously differentiable transformation function $h$ to map a complex distribution $F_{Y}$ to a simpler one $F_{Z}$ (often standard normal) (Kneib2021) via $F_{Y}=F_{Z}(h(y)).$ This allows density estimation without restrictive shape assumptions (Hothorn2018). The density $f_{Y}$ is calculated from the base density $f_{Z}$ by
@@ 2.2 Normalizing Flows and Transformation Models
NFs apply the probability-transformation theorem using a series of $K$ simple bijective transformation functions $h_{k}$ to form more expressive transformations $h(z)=h_{K}\circ h_{K-1}\circ\dots\circ h_{1}(z)$ (see, e.g. Papamakarios2021, for a comprehensive review). Other methods use a single flexible transformation like sum-of-squares polynomials (Jaini2019) or splines (Durkan2019). Conditional transformation models (CTMs) also focus on single transformation function, which, however, can be easily interpreted (see, e.g., Hothorn2018). This idea has also recently been combined with neural-networks (Baumann2021; Sick2021; Arpogaus2023; Kook2024).
A fundamental difference between NFs and CTMs is the definition of the transformation direction (Kook2024): NFs transform a simple base density $f_{Z}(z)$ into a more complex target density $f_{Y}(y)$ through a series of transformations. CTMs do it vice-versa, using only a single transformation function, often utilizing flexible Bernstein polynomials (see Section A.1 for details).
CTMs allow interpretability through the use of structured additive predictors (SAPs) (Klein2022; Rugamer2023) (see Sections A.5 and B.3). The black-box character of neural networks in NFs hinders interpretability, and variants using decision trees (Papastefanopoulos2025) and graphical models (Wehenkel2021a) have been explored in recent years.
@@ 2.3 Multivariate and Conditional Models
To model a potentially multivariate conditional distribution $F_{\mathbf{Y}|\mathbf{X}=\mathbf{x}}$, we define $z_{j}=h(y_{j},\bm{\theta}_{j})\text{ with }\bm{\theta}_{j}=c_{j}(\mathbf{y},\mathbf{x})$ where $h$ is a bijective transformation function for the element $y_{j}$ with parameters $\bm{\theta}_{j}$ obtained from the conditioner $c_{j}$ (Papamakarios2021). The latter can be an arbitrarily complex function that depends on the features $\mathbf{x}$ and, in the multivariate case, on some variables of the response $\mathbf{y}$. The limiting factor here is the computational tractability of the Jacobian in Equation 1.
@@
To make computations tractable and provide a model that is better to interpret, MCTMs define a triangular transformation $H(\bm{y},\bm{\Theta}(\bm{x}))=\left(h_{1}(y_{1}|\bm{\theta}_{1}),\ldots,h_{J}(y_{J}|y_{1},\ldots,y_{J-1},\bm{\theta}_{1},\dots,\bm{\theta}_{J})\right)^{\top}$. This Tranformation is characterized by a set of linearly combined marginal basis transformations $\tilde{h}_{j}(y_{j}|\bm{\theta}_{j})$, scaled with a $(J\times J)$ triangular coefficient matrix $\bm{\Lambda}$ to encode structural information: $h_{j}(y_{j}|y_{1},\ldots,y_{j-1},\bm{\theta}_{1},\dots,\bm{\theta}_{j})=\lambda_{j,1}\tilde{h}_{1}(y_{1}|\bm{\theta}_{1})+\ldots+\lambda_{j,j-1}\tilde{h}_{j-1}(y_{j-1}|\bm{\theta}_{j-1})+\tilde{h}_{j}(y_{j}|\bm{\theta}_{j})$. The function $\bm{\Theta}:\bm{x}\mapsto(\bm{\Lambda},\bm{\theta}_{1},\ldots,\bm{\theta}_{J})$ accounts for feature effects on the parameters (Klein2022). Hence, MCTMs can only model linear dependencies between the response variables $y_{j}$, making them an inadequate choice for complex joint distributions as shown in Figure 1. This limitation motivates the development of more flexible approaches, such as the hybrid model proposed in this paper.
@@
In the NF literature, various approaches are used to deal with computations involving the aforementioned Jacobian. Among these, autoregressive models are one of the most prevalent ones. Due to the autoregressive structure, these NFs yield a triangular Jacobian and thus the determinant simplifies to the product of the Jacobian’s diagonal entries (Papamakarios2021; Kobyzev2021). These models factorize a multivariate distribution based on the chain rule of probability. They are typically implemented by conditioning the transformation $h_{j}$ of the $j$th element of $\mathbf{y}$ on all previous elements $\mathbf{y}_{<j}=(y_{1},\dots,y_{j-1})$ or a subset of those. This can be seen as a non-linear generalization of the triangular coefficient matrix $\Lambda$ used in MCTMs (Kobyzev2021). There are two widely adopted neural-network architectures for implementing an autoregressive conditioner $c_{j}$s (Papamakarios2018) and CFs (Dinh2017).
MAFs generalize this idea by implementing the conditioner with neural networks that incorporate autoregressive constraints through parameter masking, inspired by the masked autoencoder for distribution estimation (MADE) architecture (Germain2015). The main advantage is that the parameters can be obtained in one neural network pass $\left(\bm{\theta}_{1},\ldots,\bm{\theta}_{J}\right)^{\top}=\left(c_{1}(),c_{2}(y_{1}),\ldots,c_{J}(\mathbf{y}_{<j})\right)^{\top}$. Further, if $h$ and the conditioning network are expressive enough, they remain universal approximators to transform between any two distributions (Papamakarios2021). In practice, however, the results highly depend on the ordering of the input variables. Because some orderings can be extremely difficult to learn, multiple MAFs layers are often stacked with permutations between.
@@
Graphical Transformation Models
are an extension of MCTMs, closely related to the methodology proposed in this paper. Herp2025 replaced the Gaussian copula by parameterizing the $\lambda_{i,j}$ entries of $\bm{\Lambda}$ using B-splines, dependent on the previous variables, resulting in an additive CF: $h_{j}(y_{j}|y_{1},\ldots,y_{j-1},\bm{\theta}_{1},\dots,\bm{\theta}_{j})=\lambda_{j,1}\left(\tilde{h}_{1}(y_{1}|\bm{\theta}_{1})\right)+\ldots+\lambda_{j,j-1}\left(\tilde{h}_{j-1}(y_{j-1}|\bm{\theta}_{j-1})\right)+\tilde{h}_{j}(y_{j}|\bm{\theta}_{j})$. To enhance expressiveness of the dependencies while preserving marginals, multiple CFs are chained via permutations. P-spline penalties mediate between the flexible CFs and the baseline MCTM. This framework improves interpretability of nonlinear conditional dependencies through local conditional pseudo-correlations and enables sparse undirected graphical models via a LASSO penalty promoting conditional independence. However, it offers less flexibility as a NF for decorrelation than the approach proposed here and has not yet been extended to the conditional setting.
@@ 3.1 Model Specification
This involves optimizing the parameters of the marginal transformation functions $H_{1}$ and the parameters of the autoregressive flow $H_{2}$, which are controlled by masked neural networks. Optimization can be done using any suitable gradient-based optimization algorithm, such as Adam (Kingma2017a). Once the model is trained, we can sample from the joint distribution of $\mathbf{Y}|\mathbf{X}$ by:
@@ 3.2 Interpretability
Our hybrid approach combines CTM’s interpretability with the flexibility of autoregressive NFs: The marginal transformation step ($H_{1}$) may use shifted Bernstein polynomials, which allows quantifying feature effects on the marginal distribution, similar to coefficients in linear models or generalized additive models (GAMs) (Wood2017). For instance, with a logistic base distribution, linear effect coefficients in the SAP can be interpreted as linear changes on the level of log-odds ratios (see Sections A.5 and B.3 for more details). Our model prioritizes marginal interpretability while accepting a trade-off for the dependence structure to gain flexibility via the masked autoregressive flow in step two ($H_{2}$). It is suitable for scenarios requiring understanding individual feature effects while modeling complex relationships between response variables. Understanding how features affect the dependence structure in this step is challenging and presents another open research question.
@@ 4 Numerical Experiments
Next, we evaluate our method using simulated and real-world datasets. All models are implemented using TensorFlow (v2.15.1) and TensorFlow Probability (v0.23.0) and trained using Adam (Tensorflow; TensorflowProbability; Kingma2017a). In our experiments, hyperparameters were chosen based on our previous experience and settings suggested in related works. However, we did some automated hyperparameter tuning (using Optuna) on some models applied to the simulated 2D data with performance monitored on a held-out validation set. This was mainly done to understand the influence of certain parameters, but it was not necessary to obtain a good model for each data set. We found that high-order Bernstein polynomials are needed for some datasets, suggesting the fixed knot placement of Bernstein polynomials could limit their flexibility(see Hothorn2018, for an ablation to the sensitivity of the Bernstein order).
@@ 4.1 Benchmark Datasets
We evaluate our method on five common benchmark datasets POWER, GAS, HEPMASS, MINIBOONE, and BSDS300 (see Appendix Section B.5). We follow the preprocessing steps as in Papamakarios2018 and compare a MAF and our proposed hybrid MAF (HMAF). Both models utilize masked autoregressive transformations. Each MAF layer uses a masked affine transformation, parameterized by a masked neural network followed by an invertible linear transformation (1x1 convolution) initialized with a random (non-trainable) permutation. For the HMAF, we use a marginal transformation step with Bernstein polynomials before the MAFs layers.
For both MAFs and HMAF, we employ Rational Quadratic Splines (RQS) as the transformation functions within the MAFs layers of $H_{2}$. Hyperparameters, including the number of MAFs layers, the number of hidden units in the MADE networks, and the number of bins for the RQS transformations, were chosen based on the values reported in Durkan2019 and are detailed in the supplementary material and the GitHub repository. All models were trained using the Adam optimizer with early stopping after $50$ epochs without improvements and a cosine learning rate schedule.
@@ 4.2 Simulated Data
We start with the classical bivariate (i.e. $J=2$) NFs datasets, moons and circles, exhibiting non-linear dependencies. Each dataset has $16,384$ data points (we reserved 25% for validation) generated using scikit-learn (scikit-learn). A binary feature $x$ based on spatial location is introduced. For moons, $x=1$ indicates the lower-right moon, and $x=0$ the upper-left. For circles, $x=1$ denotes the inner circle, and $x=0$ the outer. This assesses the models’ ability to capture $f(\mathbf{y}|x)$. We compare the following models:
@@
  • multivariate conditional transformation model (MCTM): Employs Bernstein polynomials of order $M=300$ for marginal transformations and a triangular matrix $\Lambda$ for the dependency structure. In the conditional case, the marginal shift parameter $\beta_{x}$ and $\Lambda_{x}$ are obtained using Bernstein polynomials of order 6, as in Klein2022. While this approach offers interpretability through SAP, the linear dependence structure constrains its capacity.
@@
All models use the Adam optimizer (Kingma2017a) with early stopping and cosine learning rate decay (Loshchilov2017). All chosen hyperparameters are reported in the supplementary material.
@@ 4.3 Malnutrition Data
Finally, we evaluate our approach using a real-world dataset on childhood malnutrition in India, with three anthropometric indices (stunting, wasting, and underweight) as response variables. The goal is to model the joint distribution conditional on the child’s age (cage). We follow the data preprocessing steps from Klein2022.
@@
  • MCTM: Identical to the model specification in Klein2022, this model combines marginal transformations with a triangular matrix $\Lambda$. We capture the influence of cage on both the marginals (using linear shifts) and on the elements of $\Lambda$ using Bernstein polynomials of order $M=6$, i.e. the marginal shift is modeled as Bernstein polynomial $\bm{\beta}_{x}=\alpha(\text{cage})^{\top}\bm{\vartheta}$, where $\alpha$ represents the Bernstein basis, and $\bm{\vartheta}$ are the coefficients.